# Physics Prospects with MeV Neutrino-Argon Charged Current Interactions using Enhanced Photon Detection in Future LArTPCs

### v1

@@ VI Summary and Outlook
We studied MeV $\nu_{e}$-Ar CC interactions and simulated the energy deposit into charge and light signals. At the generator level, an energy smearing effect is observed from hadron emission whose associated binding energy is undetectable. The dominant energy smearing at the energy deposition stage in LAr comes from neutron captures on Ar nuclei. Therefore for low energy physics it is important to tag hadron multiplicities. Further hadron PID, especially the tagging of neutrons as they are the primary knockout hadrons from these interactions according to MARLEY simulation, can help precisely correct for the energy smearing effect from neutron. The identification of protons and $\alpha$s is possible by looking for highly localized collected charge since these hadrons have high dE/dx and are expected to only span $\sim$ 1 cm in the charge readout views in LArTPCs. It is promising that neutron captures can be tagged by a combined analysis of signals from both the charge and light detection systems in LArTPC through blips analysis and timing information [33, 56]. Developing techniques for tagging MeV neutrons and delayed neutron captures on Ar nucleus at smaller scale LArTPCs utilizing information from both the charge and the light detection systems will be interesting.
An energy reconstruction based on light only calorimetry, or a combine calorimetry assuming an optimistic 75 keV charge detection threshold and average light yield of 180 PE/MeV, both show excellent energy resolution when compared to charge-only energy reconstruction. In particular, for tens of MeV $\nu_{e}$-Ar CC events, we found the energy reconstructed based on light-only calorimetry has a better resolution than combined charge and light calorimetry when hadron emission events are included. The improved energy resolution from the combined charge and light calorimetry boosts the DSNB discovery potential compared to an optimistic charge based reconstruction at the same exposure without taking into account systematic uncertainties. Once the hadron emission events are tagged, the combined calorimetry offers the best achievable DSNB sensitivity.
@@
The excellent light collection capability from extensive coverage and nanosecond timing resolution will also reduce solar neutrino analysis thresholds in a LArTPC by requiring a coincidence of pulses expected in a golden channel in the daughter ${}^{40}\mathrm{K}^{\ast}$ deexcitation leading to very low background solar neutrino samples. The background rejection using coincident light signals should be tested with a robust model of photon simulation in a LArTPC. Furthermore, a better understanding of the scintillation reconstruction is needed, including vertex resolution from light topology and timing. Reconstruction of the delayed $\gamma$ pulse also requires a precise knowledge of the scintillation time profile (with or without Xe deoping). The lifetime of the 40mK state, $480$ ns, is short compared to the few-$\mu$s primary pulse. Therefore pulse-shape discrimination will be critical for identifying the two-pulse structure. How this affects energy and vertex resolution must also be studied.
Throughout this work, the nonuniformity of the light yield in the photon detection system is not considered, but is expected to affect both physics analyses. Nevertheless, this study provides guidance to the realization of an enhanced photon detection system in future LArTPCs with high and uniform light yield. Furthermore, this work will motivate the development of more sophisticated techniques and analyses utilizing both the charge and the light detection systems in future LArTPCs to realize these low energy physics prospects.
@@ Appendix A Energy Resolution for MeV $\nu_{e}$-Ar CC Events without Hadron Emission
Appendix B Energy Smearing Matrices

### v2

@@ VI Summary and Outlook
We studied MeV $\nu_{e}$-Ar CC interactions and simulated the energy deposit into charge and light signals. At the generator level, an energy smearing effect is observed from hadron emission whose associated binding energy is undetectable. The dominant energy smearing at the energy deposition stage in LAr comes from neutron captures on Ar nuclei. Therefore for low energy physics it is important to tag hadron multiplicities. Further hadron PID, especially the tagging of neutrons as they are the primary knockout hadrons from these interactions according to MARLEY simulation, can help precisely correct for the energy smearing effect from neutrons. Developing techniques for tagging MeV neutrons and delayed neutron captures on Ar nucleus at smaller scale LArTPCs utilizing information from both the charge and the light detection systems will be interesting.
An energy reconstruction based on light only calorimetry, or a combined calorimetry assuming an optimistic 75 keV charge detection threshold and $\overline{\text{LY}}$ of 180 PE/MeV, both show excellent energy resolution when compared to charge-only energy reconstruction. In particular, for tens of MeV $\nu_{e}$-Ar CC events, we found the energy reconstructed based on light-only calorimetry has a better resolution than combined charge and light calorimetry when hadron emission events are included. The improved energy resolution from the combined charge and light calorimetry boosts the DSNB discovery potential compared to an optimistic charge based reconstruction at the same exposure without taking into account systematic uncertainties. Once the hadron emission events are tagged, the combined calorimetry offers the best achievable DSNB sensitivity.
In this study low energy radioactive background is not considered, but the related background energy deposit could be mis-clustered as a signal energy deposit. Detailed simulation with both signal and radiological background should be performed to understand event separation capability and physics sensitivity using both the charge and light detection.
@@
The excellent light collection capability from extensive coverage and nanosecond timing resolution will also reduce solar neutrino analysis thresholds in a LArTPC by requiring a coincidence of pulses expected in a golden channel in the daughter ${}^{40}\mathrm{K}^{\ast}$ deexcitation leading to very low background solar neutrino samples. The background rejection using coincident light signals should be tested with a robust model of photon simulation in a LArTPC. Furthermore, a better understanding of the scintillation reconstruction is needed, including vertex resolution from light topology and timing. Reconstruction of the delayed $\gamma$ pulse also requires a precise knowledge of the scintillation time profile (with or without Xe doping). The lifetime of the 40K state, $480$ ns, is short compared to the few-$\mu$s primary pulse. Therefore pulse-shape discrimination will be critical for identifying the two-pulse structure. How this affects energy and vertex resolution must also be studied.
The nonuniformity of light collection and its impact on energy resolution under different position reconstruction scenarios is summarized in Appendix B. This nonuniformity can be corrected once an event is located precisely by a combination of detected light signals and charge-light matching. As a result, the nonuniform LY after precise position reconstruction has little impact on energy smearing.
This study provides guidance to the realization of an enhanced photon detection system in future LArTPCs with high and uniform LY. Furthermore, this work will motivate the development of more sophisticated techniques and analyses utilizing both the charge and the light detection systems in future LArTPCs to realize these low energy physics prospects.
@@ Appendix A Energy Resolution for MeV $\nu_{e}$-Ar CC Events without Hadron Emission
Here the energy resolution is impressively low because it only considers events without any hadron emission. All hadron tagging related assumptions in the main text such as the neutron capture don’t affect this result. In addition, we only considered statistical term in the light only energy resolution. The typical contributions related to light calorimetry calibration and dark noise are neglected. For the charge side, the resolution contribution primarily comes from the applied detection threshold (cf. Fig 3 top plot) and recombination. Effects such as charge cluster and reconstruction are neglected.
@@
Appendix B Nonuniform LY Impacts
LY is not uniform across the detector volume due to effects such as optical coverage, photon propagation, and detector efficiency variations. A Geant4 simulation of the light generation and propagation in the proposed APEX design (Fig. 16) provides the LY values at different positions in the detector with an average LY of 180 PE/MeV and minimum LY of 109 PE/MeV. If an event can be precisely located in this detector, the LY nonuniformity can be corrected based on the calibrated LY map. However, location uncertainties arise when using solely the photon detection system due to the photodetector granularity. This position uncertainty introduces a mismatch between the LY for event calorimetric reconstruction and the LY for the initial generation of photons.
Figure 16: Map excerpted from Ref. [26] showing the expected LY in the central (x, y) transverse plane at z = 0 for an extended photon detection system instrumented on the field cage. Dimmer regions are present near the anode planes (with no photon detectors) and at mid-height (near the non-instrumented cathode plane).
We compare three scenarios considering the treatment of LY and its nonuniformity during the initial photon signal generation and the final reconstruction of deposited energy in light. The first scenario is the basic assumption of a uniform LY of 180 PE/MeV as described in the main text. The second scenario considers position-dependent LY following the map in Fig. 16 when generating photons and reconstructing the detected light. In this case, we assume the event position can be precisely reconstructed so that the detected energy in light can be simulated using the same LY that’s used for the initial photon generation. The precise location of an MeV event is possible after a proper and efficient charge-light (Q-L) matching. In this scenario, Eq. 5 for initial photon generation now becomes
[…]
where LY(x, y) represents the LY used to generate photons for an energy deposit at a location (x, y). And Eq. 7 for reconstructed energy deposit in light becomes
${L}_{\text{LY}}=\frac{{N}_{\text{PE}}\times 19.5\ \text{eV}}{\text{PCE}_{\text{Q-L\ matching}}}$ (13)
where $\text{PCE}_{\text{Q-L\ matching}}$ represents the photon collection efficiency assuming a precise event position can be reconstructed after Q-L matching and
[…]
Figure 17: Reconstructed total available neutrino energy for true $E_{\nu}$ = 35 MeV with different treatment of LY and its nonuniformity: a uniform LY of 180 PE/MeV as in the main text (blue); precise position reconstruction with charge-light matching (red); light-only position reconstruction (green). Top: analysis B light-only calorimetry. Middle: analysis B combined calorimetry. Bottom: analysis C combined calorimetry.
In the last scenario, we consider both the position-dependent LY and reconstructed position uncertainties using solely light detectors. This represents the worst case and we expect to see the impact from the nonuniform light collection. Here the LY for initial photon generation is the same as the second scenario in Eq. 12.
Figure 18: Reconstructed energy resolution $\sigma_{G}/\bar{E}_{G}$ for events without hadron emission reported in Appendix A with different treatment of LY and its nonuniformity: a uniform LY of 180 PE/MeV as in the main text (blue); precise position reconstruction with charge-light matching (red); light-only position reconstruction (green). Top: analysis B light-only calorimetry. Middle: analysis B combined calorimetry. Bottom: analysis C combined calorimetry.
Figure 19: Reconstructed energy resolution $\sigma_{h}/\bar{E}_{h}$ for all events reported in the main text with different treatment of LY and its nonuniformity: a uniform LY of 180 PE/MeV as in the main text (blue); precise position reconstruction with charge-light matching (red); light-only position reconstruction (green). Top: analysis B light-only calorimetry. Middle: analysis B combined calorimetry. Bottom: analysis C combined calorimetry.
However, the LY used to calculate detected energy in light is:
${L}_{\text{LY}}=\frac{{N}_{\text{PE}}\times 19.5\ \text{eV}}{\text{PCE}_{\text{Light only reconstruction}}}$ (15)
where
$\text{PCE}_{\text{Light only reconstruction}}=\frac{\text{LY}(x_{s},y_{s})}{21622}.$ (16)
Here $\text{LY}(x_{s},y_{s})$ represents a LY sampled at a reconstructed position approximated as a Gaussian smeared true event position: $x_{s}\sim N(x_{\text{true}},\ \sigma_{x}),\quad y_{s}\sim N(y_{\text{true}},\ \sigma_{y})$, where resolution $\sigma_{x}$ and $\sigma_{y}$ are the reconstructed position resolution. We take $\sigma_{x}=28$ cm and $\sigma_{y}=47$ cm for 4 MeV point energy deposit source using solely the APEX photon detectors for position reconstruction from Ref. [57]. These uncertainties are conservative since we look at true $E_{\nu}$ over 5 MeV.
The reconstructed total available energy distributions under the three scenarios of LY treatment for 35 MeV $\nu$-Ar CC events are shown in Fig. 17. The main text analysis B light-only calorimetry (Fig. 17 top), analysis B combined charge and light calorimetry (Fig. 17 middle), and analysis C combined charge and light calorimetry with hadron PID for binding energy corrections (Fig. 17 bottom) are plotted separately. The distribution with the light-only position reconstruction uncertainty (green curve) are much broader than the other two scenarios. A precise position reconstruction assumed possible via charge-light matching (red) removes the effect on position dependendt LY and result in a similar performance to the uniform LY scenario (blue).
Fig. 18 and Fig. 19 present energy resolutions $\sigma_{G}/\bar{E}_{G}$ and $\sigma_{h}/\bar{E}_{h}$ defined in the main text for neutrino energy from 5 to 50 MeV. The resolution with nonuniform LY and the uncertainty of the position of the event vertex has a larger impact on light-only calorimetric reconstruction than the combined calorimetric reconstruction.
Lastly, we comment on the precise position reconstruction scenario using charge-light matching (red curves in all plots) where the effect LY nonuniformity is removed. With an enhanced light detection system such as APEX, the light-only reconstructed position resolution is on the order of tens of cm. We expect once the interesting activity is located to this tens-of-cm region solely by the light detector, we can perform more efficient and accurate charge-light matching to pin point the activity more precisely, on the order of 1 cm (i.e., LArTPC charge readout position resolution). This helps us to correct for the LY nonuniformity effects.
Appendix C Energy Smearing Matrices